# Active Sensing for Multiuser Beam Tracking with Reconfigurable Intelligent Surface

### v2

@@ IV-C1 RIS Sensing Vector Design
In the pilot stage of the $(t-1)$-th frame, the LSTM cell state $\bm{c}_{k}^{(t-1)}$’s that encapsulate the temporal channel correlations for the $K$ mobile UEs based on the previous frames are fed into a GNN. The GNN then outputs the RIS sensing vectors $\{\bm{v}^{(t)}_{\ell}\}^{L}_{\ell=1}$ for the next round of CSI acquisition (i.e., pilot stage of the $t$-th frame). We observe empirically that it is also possible to replace the LSTM cell state $\bm{c}_{k}^{(t-1)}$ with the hidden state $\bm{h}_{k}^{(t-1)}$ as the GNN input.
@@ IV-C2 State Vector Update
In the $t$-th frame, $K$ LSTM cells with shared weights take in the most recent channel observation $\bar{\bm{Y}}_{k}^{(t)}$’s, alongside the preceding cell state $\bm{c}_{k}^{(t-1)}$’s and hidden state $\bm{h}_{k}^{(t-1)}$’s, to generate the new cell states $\bm{c}_{k}^{(t)}$’s and hidden state $\bm{h}_{k}^{(t)}$’s. By convention, we initialize $\bm{c}_{k}^{(-1)}=\bm{h}_{k}^{(-1)}=\bm{0}$, and $\bar{\bm{Y}}_{k}^{(0)}=\bm{1}$ for all $k$ in $\{0,\dots,K\}$.
@@ IV-C3 Downlink Beam Alignment
Utilizing the updated LSTM cell state $\bm{c}_{k}^{(t)}$’s as input, the proposed GNN computes the downlink RIS reflection coefficients $\bm{w}^{(t)}$ and power allocations $\{\bm{p}^{(t)},\bm{\lambda}^{(t)}\}$. The AP then estimates the effective low-dimensional channel $\bm{h}_{c,k}^{(t,0)}$’s based on the additional $\tau_{w}$ pilot symbols transmitted via the reflection of the $(L+1)$-th RIS sensing vector $\bm{v}_{L+1}^{(t)}=\bm{w}^{(t)}$. With the estimated low-dimensional channel $\check{\bm{h}}_{c,k}^{(t,0)}$ and designed power allocations $\{\bm{p}^{(t)},\bm{\lambda}^{(t)}\}$, the downlink beamforming vector $\bm{b}^{(t)}_{k}$’s for subsequent data transmissions are computed via (9).
@@ IV-D Neural Network Training and Inference
In the simulations, we observe that it is possible to further improve the performance of directly using the power allocations $\{\bm{p}^{(t)},\bm{\lambda}^{(t)}\}$ designed by neural network. This is primarily because the proposed active sensing framework only has moderate size of the hidden neurons, which restricts its expressive capacity. It is possible to adopt a refined approach using the fact that the effective low-dimensional channel $\bm{h}_{c,k}^{(t,0)}$’s can be easily estimated according to (11). By employing a fixed-point iterations method [38, 39] which takes the estimated $\check{\bm{h}}_{c,k}^{(t,0)}$’s as input, we can analytically determine the optimal power allocations. This method allows for a more precise adjustment of power allocations, ensuring that the beamforming vectors computed therefrom are better optimized.
More specifically, the optimal power allocation $\bm{\lambda}^{(t)}_{*}$ is the unique positive solution of the following fixed-point equation:
@@
$\bm{p}^{(t)}_{*}=(\bm{I}_{K}-\tau^{(t)}_{*}\bm{D}^{(t)}\bm{F}^{(t)})^{-1}\tau^{(t)}_{*}\sigma_{d}^{2}\bm{D}^{(t)}\bm{1}_{K}\ ,$ (17)
where $\bm{D}^{(t)}=\text{diag}(\frac{1}{|(\check{\bm{h}}_{c,1}^{(t,0)})^{\sf{H}}\widetilde{\bm{b}}^{(t)}_{1}|^{2}},\dots,\frac{1}{|(\check{\bm{h}}_{c,K}^{(t,0)})^{\sf{H}}\widetilde{\bm{b}}^{(t)}_{K}|^{2}})$ and $[\bm{F}^{(t)}]_{k,i}=1/|(\check{\bm{h}}_{c,k}^{(t,0)})^{\sf{H}}\widetilde{\bm{b}}^{(t)}_{i}|^{2}$, if $k\neq i$ and 00 otherwise. Here, $\widetilde{\bm{b}}^{(t)}_{k}$’s denote the beamforming directions which can be computed based on $\bm{\lambda}^{(t)}_{*}$ and $\check{\bm{h}}_{c,k}^{(t,0)}$’s via (9).
@@ V-A2 Channel Model
where the correlation coefficient $\rho=0.9995$, perturbation term $\mu^{(k,n)}\sim\mathcal{CN}(0,1)$, and $[\widetilde{\bm{h}}_{d,k}^{(1,1)}]_{i}\sim\mathcal{CN}(0,1)$. We assume that the RIS is deployed at a location where multiple paths exist between the AP and the RIS, and a LOS channel exists between the RIS and each mobile UE. More specifically, the AP-RIS link $\bm{G}^{(t,n)}$ and the RIS-UE link $\bm{h}_{r,k}^{(t,n)}$’s follow the Rician fading model as:
$\bm{G}^{(t,n)}=\beta_{G}^{(t,n)}(\sqrt{\varepsilon/(1+\varepsilon)}\ \sum_{i}^{N_{G}}\widehat{\bm{G}}_{i}+\sqrt{1/(1+\varepsilon)}\ \widetilde{\bm{G}}^{(t,n)}),$ (21)
$\bm{h}_{r,k}^{(t,n)}=\beta_{r,k}^{(t,n)}(\sqrt{\varepsilon/(1+\varepsilon)}\ \widehat{\bm{h}}_{r,k}^{(t,n)}+\sqrt{1/(1+\varepsilon)}\ \widetilde{\bm{h}}_{r,k}^{(t,n)}),$ (22)
where $N_{G}$ is the number of paths between the AP and RIS, and $\varepsilon$ is the Rician factor which is set to be $10$. The path loss (in dB) of the AP-RIS link $\beta_{G}$ and the RIS-UE link $\beta_{r}^{(t,n)}$ are modeled as $28+16.9\log(l_{G})$ and $28+16.9\log(l_{r}^{(t,n)})$, respectively, where $l_{G}$ and $l_{r}^{(t,n)}$ are the distance of the corresponding links. Similar to (20), the distribution of the entries of the NLOS components of the AP-RIS link $[\widetilde{\bm{G}}^{(t,n)}]_{ij}$ and the RIS-UE link $[\widetilde{\bm{h}}_{r,k}^{(t,n)}]_{i}$’s are modeled by the same stationary Gauss-Markov process, with $[\widetilde{\bm{G}}^{(1,1)}]_{ij}\sim\mathcal{CN}(0,1)$ and $[\widetilde{\bm{h}}_{r,k}^{(1,1)}]_{i}\sim\mathcal{CN}(0,1)$.
@@
$=e^{j\frac{2\pi d_{\text{RIS}}}{\lambda_{c}}\{i_{1}(u,v)\sin(\phi_{3,k}^{(t,n)})\cos(\theta_{3,k}^{(t,n)})+i_{2}(u,v)\sin(\theta_{3,k}^{(t,n)})\}}.$
Here, $\lambda_{c}$ is the carrier wavelength, $d_{\text{RIS}}$ is the distance between two adjacent RIS elements, $i_{1}(u,v)=\bmod(u-1,v)$, and $i_{2}(u,v)=\lfloor\frac{u-1}{v}\rfloor$, where $v$ is the number of columns of the RIS. Without loss of generality, we set $d_{\text{RIS}}=\lambda_{c}/2$ and $v=10$ in simulations. Similarly, let $\phi_{2,i}$ and $\phi_{1,i}$ denote the $i$-th azimuth angle of departure (AoD) and azimuth AoA from the RIS to the AP, corresponding to the $i$-th path $\widehat{\bm{G}}_{i}$ between the AP and the RIS. Then we can write $\widehat{\bm{G}}_{i}=\bm{a}_{\text{AP}}(\phi_{1,i})\bm{a}_{\text{RIS}}(\theta_{2,i},\phi_{2,i})^{\sf{H}}$, where $\theta_{2,i}=0$ as shown in Fig. 8 and the steering vector of the AP is given by:
@@ V-B Implementation Details
The proposed active sensing framework is implemented on Tensorflow [40]. For the LSTM cells, the dimensions of hidden state $\bm{h}_{k}^{(t)}$’s, cell state $\bm{c}_{k}^{(t)}$’s, and hidden layers $\bm{u}_{c}$, $\bm{u}_{q}$, $\bm{u}_{i}$, $\bm{u}_{o}$, $\bm{w}_{c}$, $\bm{w}_{q}$, $\bm{w}_{i}$, $\bm{w}_{o}$ are set to be $512$. We adopt a GNN with $D=2$ updating layers. Within the GNN architecture, the DNNs $f_{1}(\cdot)$, $f_{2}(\cdot)$, and $f_{3}(\cdot)$ are of size $512\times 1024\times 512$, and the size of DNNs $f_{4}(\cdot)$, $f_{5}(\cdot)$, $f_{6}(\cdot)$, $f_{7}(\cdot)$, $f_{8}(\cdot)$, $f_{9}(\cdot)$, $f_{w}(\cdot)$, $f_{V}(\cdot)$ are $512\times 512\times 512$. We concatenate $U=20$ active sensing units to train the proposed framework. In each training epoch, the neural network samples $6400$ training data including random channel and noise realizations. We terminate the training process if the validation loss does not decrease over $30$ consecutive training epochs. An instance of trajectories for $K=3$ mobile UEs is shown in Fig. 8.
We assume there are $N_{G}=8$ paths between the AP and RIS. Each transmission frame contains $N=3$ blocks. In the pilot stage (the 00-th block) of each frame, there are $4$ sub-blocks, i.e., $L=3$. The effective downlink channel $\bm{h}_{c,k}^{(t,0)}$’s are estimated in the $4$-th sub-block with $\tau_{w}=10$ pilots.
@@ V-C Benchmark Methods
Given the perfect CSI of the channel $\bm{G}^{(t,1)}$, $\bm{h}_{\mathrm{d},k}^{(t,1)}$’s, and $\bm{h}_{\mathrm{r},k}^{(t,1)}$’s, we can find a performance benchmark for the proposed active sensing framework by solving the following optimization problem:
@@
subject to $\sum_{k}\|\bm{b}_{k}\|_{2}^{2}\leq P_{d}$ and $|[\bm{w}^{(t)}]_{n}|=1$ using the block coordinate descent (BCD) algorithm, which iteratively optimizes the beamforming matrix $\bm{B}^{(t)}$ and the downlink RIS reflection coefficients $\bm{w}^{(t)}$. Specifically, in each iteration, we optimize $\bm{B}^{(t)}$ using the fixed-point iterations (16) - (17) with $\bm{w}^{(t)}$ fixed [38], and then optimize $\bm{w}^{(t)}$ using the Riemannian conjugate gradient (RCG) method with fixed $\bm{B}^{(t)}$ [41, 42].
@@
In the first $L$ sub-blocks of each transmission frame, the AP receives pilots through a fixed set of the sensing vectors $\{\bm{v}_{\ell}\}^{L}_{\ell=1}$ which are generated either randomly or learned from channel statistics. The GNN as described in the proposed approach is employed to map the received pilots $\bar{\bm{Y}}_{k}^{(t)}$’s to the downlink RIS reflection coefficients $\bm{w}^{(t)}$. Subsequently, $\tau_{w}$ additional pilot symbols are transmitted to estimate the effective channel $\bm{h}_{c,k}^{(t,0)}$’s for power allocation refinement. The sensing vector $\bm{v}_{L+1}^{(t)}$ is set to be $\bm{w}^{(t)}$ in this refinement stage.
@@
This baseline is similar to the proposed approach, except that the RIS sensing vectors $\{\bm{v}^{(t)}_{\ell}\}^{L}_{\ell=1}$ are not actively designed across different transmission frames. Without active sensing, $\bm{v}^{(t)}_{\ell}$’s can be designed based on the following three schemes: (i) the phases of each $\bm{v}^{(t)}_{\ell}$ are randomly drawn from a uniform distribution over $[0,2\pi)$ [24]; (ii) $\bm{v}^{(t)}_{\ell}$’s are learned from statistics according to channel realizations in the training phase [6]; (iii) using the current RIS reflection coefficients $\bm{w}^{(t)}$ for the next round of pilot acquisition [18], i.e., $\bm{v}_{\ell}^{(t+1)}=\bm{w}^{(t)}$, $\ell=1,\dots,L$. In (i), (ii), and (iii), $\tau_{w}$ additional pilot symbols are transmitted in the $(L+1)$-th sub-block with $\bm{v}_{L+1}^{(t)}=\bm{w}^{(t)}$ to estimate the effective channel $\bm{h}_{c,k}^{(t,0)}$’s for power allocation refinement.
@@
Figure 9: Average minimum user rate at the $1$-st block of each frame (i.e., $\min_{k}\{R_{k}^{(t,1)}\}$) for different beam tracking methods. Here, $L=3$, $M=8$, $N_{r}=100$, $K=3$, $N_{G}=8$, $\tau_{w}=10$.
Figure 10: The impact of different number of sensing vectors $L$’s within each frame on different beam tracking methods. Here, $M=8$, $N_{r}=100$, $K=3$, $N_{G}=8$, $\tau_{w}=10$.
V-D Performance Comparison

### v3

@@ IV-C1 RIS Sensing Vector Design
In the pilot stage of the $(t-1)$-th frame, the LSTM cell states $\bm{c}_{k}^{(t-1)}$ that encapsulate the temporal channel correlations for the $K$ mobile UEs based on the previous frames are fed into a GNN. The GNN then outputs the RIS sensing vectors $\{\bm{v}^{(t)}_{\ell}\}^{L}_{\ell=1}$ for the next round of CSI acquisition (i.e., pilot stage of the $t$-th frame). We observe empirically that it is also possible to replace the LSTM cell state $\bm{c}_{k}^{(t-1)}$ with the hidden state $\bm{h}_{k}^{(t-1)}$ as the GNN input.
@@ IV-C2 State Vector Update
In the $t$-th frame, $K$ LSTM cells with shared weights take in the most recent channel observations $\bar{\bm{Y}}_{k}^{(t)}$, alongside the preceding cell states $\bm{c}_{k}^{(t-1)}$ and hidden states $\bm{h}_{k}^{(t-1)}$, to generate the new cell states $\bm{c}_{k}^{(t)}$ and hidden states $\bm{h}_{k}^{(t)}$. By convention, we initialize $\bm{c}_{k}^{(-1)}=\bm{h}_{k}^{(-1)}=\bm{0}$, and $\bar{\bm{Y}}_{k}^{(0)}=\bm{1}$ for all $k$ in $\{0,\dots,K\}$.
@@ IV-C3 Downlink Beam Alignment
Utilizing the updated LSTM cell states $\bm{c}_{k}^{(t)}$ as input, the proposed GNN computes the downlink RIS reflection coefficients $\bm{w}^{(t)}$ and power allocations $\{\bm{p}^{(t)},\bm{\lambda}^{(t)}\}$. The AP then estimates the effective low-dimensional channels $\bm{h}_{c,k}^{(t,0)}$ based on the additional $\tau_{w}$ pilot symbols transmitted via the reflection of the $(L+1)$-th RIS sensing vector $\bm{v}_{L+1}^{(t)}=\bm{w}^{(t)}$. With the estimated low-dimensional channel $\check{\bm{h}}_{c,k}^{(t,0)}$ and designed power allocations $\{\bm{p}^{(t)},\bm{\lambda}^{(t)}\}$, the downlink beamforming vectors $\bm{b}^{(t)}_{k}$ for subsequent data transmissions are computed via (9).
@@ IV-D Neural Network Training and Inference
In the simulations, we observe that it is possible to further improve the performance of directly using the power allocations $\{\bm{p}^{(t)},\bm{\lambda}^{(t)}\}$ designed by neural network. This is primarily because the proposed active sensing framework only has moderate size of the hidden neurons, which restricts its expressive capacity. It is possible to adopt a refined approach using the fact that the effective low-dimensional channels $\bm{h}_{c,k}^{(t,0)}$ can be easily estimated according to (11). By employing a fixed-point iterations method [40, 41] which takes the estimated $\check{\bm{h}}_{c,k}^{(t,0)}$’s as input, we can analytically determine the optimal power allocations. This method allows for a more precise adjustment of power allocations, ensuring that the beamforming vectors computed therefrom are better optimized.
More specifically, the optimal power allocation $\bm{\lambda}^{(t)}_{*}$ is the unique positive solution of the following set of fixed-point equations, for $k=1,\cdots,K$,
@@
$\bm{p}^{(t)}_{*}=\left(\bm{I}_{K}-\tau^{(t)}_{*}\bm{D}^{(t)}\bm{F}^{(t)}\right)^{-1}\tau^{(t)}_{*}\sigma_{d}^{2}\bm{D}^{(t)}\bm{1}_{K},$ (17)
where $\bm{D}^{(t)}=\text{diag}\left(\frac{1}{|(\check{\bm{h}}_{c,1}^{(t,0)})^{\sf{H}}\widetilde{\bm{b}}^{(t)}_{1}|^{2}},\dots,\frac{1}{|(\check{\bm{h}}_{c,K}^{(t,0)})^{\sf{H}}\widetilde{\bm{b}}^{(t)}_{K}|^{2}}\right)$ and $[\bm{F}^{(t)}]_{k,i}=\frac{1}{\left|\left(\check{\bm{h}}_{c,k}^{(t,0)}\right)^{\sf{H}}\widetilde{\bm{b}}^{(t)}_{i}\right|^{2}}$, if $k\neq i$ and 00 otherwise. Here, $\widetilde{\bm{b}}^{(t)}_{k}$’s denote the beamforming directions which can be computed based on $\bm{\lambda}^{(t)}_{*}$ and $\check{\bm{h}}_{c,k}^{(t,0)}$’s via (9).
@@ V-A2 Channel Model
where the correlation coefficient $\rho=0.9995$, perturbation term $\mu^{(k,n)}\sim\mathcal{CN}(0,1)$, and $[\widetilde{\bm{h}}_{d,k}^{(1,1)}]_{i}\sim\mathcal{CN}(0,1)$. We assume that the RIS is deployed at a location where multiple paths exist between the AP and the RIS, and a LOS channel exists between the RIS and each mobile UE. More specifically, the AP-RIS link $\bm{G}^{(t,n)}$ and the RIS-UE links $\bm{h}_{r,k}^{(t,n)}$ follow the Rician fading model as:
$\bm{G}^{(t,n)}\!=\!\beta_{G}^{(t,n)}\!\left(\sqrt{\varepsilon/(1+\varepsilon)}\sum_{i}^{N_{G}}\widehat{\bm{G}}_{i}+\sqrt{1/(1+\varepsilon)}\ \widetilde{\bm{G}}^{(t,n)}\right),$ (21)
$\bm{h}_{r,k}^{(t,n)}\!=\!\beta_{r,k}^{(t,n)}\!\left(\sqrt{\varepsilon/(1+\varepsilon)}\ \widehat{\bm{h}}_{r,k}^{(t,n)}+\sqrt{1/(1+\varepsilon)}\ \widetilde{\bm{h}}_{r,k}^{(t,n)}\right),$ (22)
where $N_{G}$ is the number of paths between the AP and RIS, and $\varepsilon$ is the Rician factor which is set to be $10$. The path loss (in dB) of the AP-RIS link $\beta_{G}$ and the RIS-UE link $\beta_{r}^{(t,n)}$ are modeled as $28+16.9\log(l_{G})$ and $28+16.9\log(l_{r,k}^{(t,n)})$, respectively, where $l_{G}$ and $l_{r,k}^{(t,n)}$ are the distance of the corresponding links. Similar to (20), the distribution of the entries of the NLOS components of the AP-RIS link $[\widetilde{\bm{G}}^{(t,n)}]_{ij}$ and the RIS-UE links $[\widetilde{\bm{h}}_{r,k}^{(t,n)}]_{i}$ are modeled by the same stationary Gauss-Markov process, with $[\widetilde{\bm{G}}^{(1,1)}]_{ij}\sim\mathcal{CN}(0,1)$ and $[\widetilde{\bm{h}}_{r,k}^{(1,1)}]_{i}\sim\mathcal{CN}(0,1)$.
@@
$=e^{j\frac{2\pi d_{\text{RIS}}}{\lambda_{c}}\left(i_{1}(u,v)\sin(\phi_{3,k}^{(t,n)})\cos(\theta_{3,k}^{(t,n)})+i_{2}(u,v)\sin(\theta_{3,k}^{(t,n)})\right)}.$
Here, $\lambda_{c}$ is the carrier wavelength, $d_{\text{RIS}}$ is the distance between two adjacent RIS elements, $i_{1}(u,v)=\bmod(u-1,v)$, and $i_{2}(u,v)=\lfloor\frac{u-1}{v}\rfloor$, where $v$ is the number of columns of the RIS. Without loss of generality, we set $d_{\text{RIS}}=\lambda_{c}/2$ and $v=10$ in simulations. As shown in Fig. 8, the RIS is located at $(0\text{m},0\text{m},0\text{m})$. Let $(x_{k}^{(t,n)},y_{k}^{(t,n)},z_{k})$ denote the coordinates of the UE $k$ with $z_{k}=-20\text{m}$, we have $\sin(\phi_{3,k}^{(t,n)})\cos(\theta_{3,k}^{(t,n)})=y_{k}^{(t,n)}/l_{r,k}^{(t,n)}$ and $\sin(\theta_{3,k}^{(t,n)})=z_{k}/l_{r,k}^{(t,n)}$. Similarly, let $\phi_{2,i}$ and $\phi_{1,i}$ denote the $i$-th azimuth angle of departure (AoD) and azimuth AoA from the RIS to the AP, corresponding to the $i$-th path $\widehat{\bm{G}}_{i}$ between the AP and the RIS. Then we can write $\widehat{\bm{G}}_{i}=\bm{a}_{\text{AP}}(\phi_{1,i})\bm{a}_{\text{RIS}}(\theta_{2,i},\phi_{2,i})^{\sf{H}}$, where $\theta_{2,i}=0$ as shown in Fig. 8 and the steering vector of the AP is given by:
@@ V-B Implementation Details
The proposed active sensing framework is implemented on Tensorflow [42]. For the LSTM cells, the dimensions of hidden states $\bm{h}_{k}^{(t)}$, cell states $\bm{c}_{k}^{(t)}$, and hidden layers $\bm{u}_{c}$, $\bm{u}_{q}$, $\bm{u}_{i}$, $\bm{u}_{o}$, $\bm{w}_{c}$, $\bm{w}_{q}$, $\bm{w}_{i}$, $\bm{w}_{o}$ are set to be $512$. We adopt a GNN with $D=2$ updating layers. Within the GNN architecture, the DNNs $f_{1}(\cdot)$, $f_{2}(\cdot)$, and $f_{3}(\cdot)$ are of size $512\times 1024\times 512$, and the size of DNNs $f_{4}(\cdot)$, $f_{5}(\cdot)$, $f_{6}(\cdot)$, $f_{7}(\cdot)$, $f_{8}(\cdot)$, $f_{9}(\cdot)$, $f_{w}(\cdot)$, $f_{V}(\cdot)$ are $512\times 512\times 512$. We concatenate $U=20$ active sensing units to train the proposed framework. In each training epoch, the neural network samples $6400$ training data including random channel and noise realizations. We terminate the training process if the validation loss does not decrease over $30$ consecutive training epochs. An instance of trajectories for $K=3$ mobile UEs is shown in Fig. 8.
We assume there are $N_{G}=8$ paths between the AP and RIS. Each transmission frame contains $N=3$ blocks. In the pilot stage (the 00-th block) of each frame, there are $4$ sub-blocks, i.e., $L=3$. The effective downlink channels $\bm{h}_{c,k}^{(t,0)}$ are estimated in the $4$-th sub-block with $\tau_{w}=10$ pilots.
@@ V-C Benchmark Methods
Given the perfect CSI of the channels $\bm{G}^{(t,1)}$, $\bm{h}_{\mathrm{d},k}^{(t,1)}$, and $\bm{h}_{\mathrm{r},k}^{(t,1)}$, we can find a performance benchmark for the proposed active sensing framework by solving the following optimization problem:
@@
subject to $\sum_{k}\|\bm{b}^{(t)}_{k}\|_{2}^{2}\leq P_{d}$ and $|[\bm{w}^{(t)}]_{n}|=1$ using the block coordinate descent (BCD) algorithm, which iteratively optimizes the beamforming matrix $\bm{B}^{(t)}$ and the downlink RIS reflection coefficients $\bm{w}^{(t)}$. Specifically, in each iteration, we optimize $\bm{B}^{(t)}$ using the fixed-point iterations (16) - (17) with $\bm{w}^{(t)}$ fixed [40], and then optimize $\bm{w}^{(t)}$ using the Riemannian conjugate gradient (RCG) method with fixed $\bm{B}^{(t)}$ [43, 44].
@@
In the first $L$ sub-blocks of each transmission frame, the AP receives pilots through a fixed set of the sensing vectors $\{\bm{v}_{\ell}\}^{L}_{\ell=1}$ which are generated either randomly or learned from channel statistics. The GNN as described in the proposed approach is employed to map the received pilots $\bar{\bm{Y}}_{k}^{(t)}$ to the downlink RIS reflection coefficients $\bm{w}^{(t)}$. Subsequently, $\tau_{w}$ additional pilot symbols are transmitted to estimate the effective channels $\bm{h}_{c,k}^{(t,0)}$ for power allocation refinement. The sensing vector $\bm{v}_{L+1}^{(t)}$ is set to be $\bm{w}^{(t)}$ in this refinement stage.
@@
This baseline is similar to the proposed approach, except that the RIS sensing vectors $\{\bm{v}^{(t)}_{\ell}\}^{L}_{\ell=1}$ are not actively designed across different transmission frames. Without active sensing, $\bm{v}^{(t)}_{\ell}$’s can be designed based on the following three schemes: (i) the phases of each $\bm{v}^{(t)}_{\ell}$ are randomly drawn from a uniform distribution over $[0,2\pi)$ [24]; (ii) $\bm{v}^{(t)}_{\ell}$’s are learned from statistics according to channel realizations in the training phase [6]; (iii) using the current RIS reflection coefficients $\bm{w}^{(t)}$ for the next round of pilot acquisition [18], i.e., $\bm{v}_{\ell}^{(t+1)}=\bm{w}^{(t)}$, $\ell=1,\dots,L$. In (i), (ii), and (iii), $\tau_{w}$ additional pilot symbols are transmitted in the $(L+1)$-th sub-block with $\bm{v}_{L+1}^{(t)}=\bm{w}^{(t)}$ to estimate the effective channels $\bm{h}_{c,k}^{(t,0)}$ for power allocation refinement.
GNN & LSTM with GNN Directly Designing $\bm{B}^{(t)}$
This benchmark is similar to the proposed active sensing framework, except that the beamforming structure (9) is not leveraged to design the AP beamforming matrix $\bm{B}^{(t)}$ in the training stage. In this case, instead of outputting the power allocations $\{\bm{p}^{(t)},\bm{\lambda}^{(t)}\}$, the GNN directly outputs the exact downlink beamforming vectors $\bm{b}^{(t)}_{k}$ for each user. During inference, $\tau_{w}$ additional pilot symbols are transmitted in the $(L+1)$-th sub-block with $\bm{v}_{L+1}^{(t)}=\bm{w}^{(t)}$ for power allocation refinement.
@@
Figure 9: Average minimum user rate at the $1$-st block of each frame (i.e., $\min_{k}\{R_{k}^{(t,1)}\}$) for different beam tracking methods. Here, $L=3$, $M=8$, $N_{r}=100$, $K=3$, $N_{G}=8$, $\tau_{w}=10$, $P_{u}=5\text{dBm}$, and $P_{d}=10\text{dBm}$.
V-D Performance Evaluation